# Search for neutrino emission from blazar $\gamma$-ray flares accounting for possible neutrino time delays

### v1

@@ 3 Methods
$\left(\frac{d^{\alpha}_{\nu b}}{\tilde{\alpha}_{\nu}^{\pm}}\right)^{2}+\left(\frac{d_{\nu b}^{\delta}}{\tilde{\delta}_{\nu}^{\pm}}\right)^{2}\leq 1,$ (3.1)
@@
$\tilde{\alpha}_{\nu}^{\pm}=\alpha_{\nu}^{\pm}+\Delta,$ (3.2)
$\tilde{\delta}_{\nu}^{\pm}=\delta_{\nu}^{\pm}+\Delta,$ (3.3)
where $d^{\alpha}_{\nu b}$ is the difference between blazar’s $b$ and neutrino’s $\nu$ RAs, $d^{\delta}_{\nu b}$ is the difference between blazar’s and neutrino’s declinations, the sign ++ or - in $\alpha_{\nu}^{\pm}$ and $\delta_{\nu}^{\pm}$ is chosen according to the sign of $d^{\alpha}_{\nu b}$ and $d^{\delta}_{\nu b}$; $\Delta=0.78^{\circ}$ is the assumed systematic angular uncertainty. This value was found [14] to give the optimal correlation strength between the radio-bright blazars and neutrino alerts, and we keep it fixed to avoid statistical penalising for multiple trials. A similar value of $1^{\circ}$ was found [19] to be optimal in the search for spatio-temporal correlations between the multiwavelength light curves of the CGRaBS blazars [44] and the IceCube high-energy neutrinos. In a dedicated study with synthesised catalogues and a controlled number of synthetic blazar-neutrino correlations, Ref. [46] showed that enlarging the value of $\Delta$ does not lead to an increase in chances of finding a false-positive statistically significant correlation.
Following Eq. (1) in Ref. [46], we define for each of the alerts its enlarged angular uncertainty area as
$U_{\nu}=\frac{\pi}{4}\left(\tilde{\alpha}^{+}_{\nu}\tilde{\delta}^{+}_{\nu}+\tilde{\alpha}^{-}_{\nu}\tilde{\delta}^{+}_{\nu}+\tilde{\alpha}^{-}_{\nu}\tilde{\delta}^{-}_{\nu}+\tilde{\alpha}^{+}_{\nu}\tilde{\delta}^{-}_{\nu}\right),$ (3.4)
@@
If neutrinos are produced in the external photon fields, approximate linearity between the $\gamma$ and $\nu$ fluxes may be expected. According to [47], the Doppler amplification $\propto D^{p}$ makes the observed neutrino flux boosted with $p=4$ if neutrinos are produced in $p\gamma$ interactions on internal photon fields co-moving with the jet frame. However, $p\geq 5$ is expected if $p\gamma$ interactions are dominated by the photon fields external to the relativistic jet, with especially strong amplifications $p>5$ if $p\gamma$ photopion production occurs near the threshold. The $\gamma$-ray external Compton emission is, in turn, produced with amplification $\propto p^{6}$, making $F_{\nu}\propto F_{\gamma}^{\geq 5/6}\sim F_{\gamma}$, which justifies our assumption of the linear $\gamma-\nu$ flux scaling. Ref. [48] found that neutrino-emitting blazar candidates are statistically compatible with hypotheses of both a linear correlation and no correlation between neutrino and gamma-ray energy fluxes.
We expect the jet-frame time delay $t^{\prime}_{\mathrm{delay}}$ to relate to the photopion proton energy-loss timescale $t^{\prime}_{p\gamma}$ and/or proton acceleration timescale $t^{\prime}_{\mathrm{acc}}$. They can vary significantly depending on the proton energy and parameters of photon fields serving as the target for the photopion production process, on the magnetic field, and proton acceleration mechanism. A priori, we do not know the exact value of $t^{\prime}_{\mathrm{delay}}$. Thus, we perform a scan over parameter $t^{\prime}_{\mathrm{delay}}\sim U[t^{\prime}_{\min};t^{\prime}_{\max}]$, $t^{\prime}_{\min}=\Delta_{F}=1$ month corresponding to the bin size $\Delta_{F}$ of the Fermi-LAT light curves. Ref. [27], in modelling the brightest Fermi-LAT blazar flare of 3C 454.3, obtained that close to the threshold of the photopion production around $E_{\nu}\sim 100$ TeV, $t^{\prime}_{\mathrm{delay}}\approx t^{\prime}_{p\gamma}\approx t^{\prime}_{\mathrm{acc}}\sim 10^{9}$ s (see Fig. 12 in Ref. [27]). Thus, we adopt $t^{\prime}_{\max}=3\times 10^{9}$ s $3.47\times 10^{4}$ d being agnostic of the exact mechanism of the possible $\nu$ delay and no particular preference to any delay between $t^{\prime}_{\min}$ and $t^{\prime}_{\max}$.
@@
$w_{\nu}(t^{\prime}_{\mathrm{delay}})=\sum_{a}\left[s_{\nu}\rho_{\nu}G_{\nu a}(t^{\gamma\max}_{a},t_{\nu};t^{\prime}_{\mathrm{\,delay}},D_{a},z_{a})F^{\gamma}_{a}(t^{\gamma\max}_{a})\right],$ (3.5)
where the sum $\sum_{a}$ runs over all sources within the uncertainty region of neutrino $\nu$ (as was pointed out by [49], a single neutrino cannot come from more than one source, but having several blazars in its angular uncertainty region increases the probability of this neutrino coming from a blazar), $s_{\nu}$ is the alert signalness, $\rho_{\nu}$ is the factor penalising alerts with too large uncertainties. Following recommendations of [46], we use a top-hat spatial factor penalising alerts with too large angular uncertainties:
$\rho_{\nu}=\begin{cases}1,\mathrm{\,if\,}U_{\nu}\leq\tilde{U}_{\nu}\\ \displaystyle{\left(\frac{\tilde{U}_{\nu}}{U_{\nu}}\right)},\mathrm{\,if\,}U_{\nu}>\tilde{U}_{\nu},\end{cases}$ (3.6)
where $\tilde{U}_{\nu}=15.6$ sq. deg. is the median angular uncertainty (see Eq. 3.4) of IceCat-1 alerts (after the extension by $\Delta=0.78^{\circ}$).
The key point of the current paper is the search for neutrino-blazar correlations taking into account the possibility of neutrino emission lagging behind the $\gamma$-ray flare due to slowness of the $p\gamma$ photopion production and/or proton acceleration as was suggested in [27]. Given that even during the brightest blazar flares, the IceCube neutrino yield is of the order of $10^{-3}-10^{-2}$ muon neutrinos at energies $\gtrsim 100$ TeV [50, 51, 52] and during quieter activity periods even lower (assuming proportionality between $\gamma$-ray and $\nu$ fluxes), we assume that if the neutrino is indeed associated with the blazar, it must have come in a causal connection with the major flare. Thus, the last term in Eq. (3.5) denotes the global maximum of the Fermi-LAT light curve of blazar $a$ detected at $t^{\gamma\max}_{a}$ among the times prior to the associated neutrino arrival ($\Delta_{F}=1$ month to account for the size of the time bin and still allow for instantaneous associations) normalized by $1.66\times 10^{-5}$ cm-2 s-1 (the maximum value of monthly-averaged fluxes among the studied blazars within the whole Fermi-LAT observation history observed from FSRQ 3C 454.3 during the brightest Fermi-LAT blazar flare in November 2010 [53, 54, 27]). As for the temporal term $G_{\nu a}$ accounting for the neutrino delay, we define it linking to the time $t^{\gamma\max}_{a}$ of the global maximum of the associated blazar light curve prior to the neutrino arrival as follows
@@
$t_{a\mathrm{\,delay}}(t^{\prime}_{\mathrm{delay}},D_{a},z_{a})=\max\left\{\Delta_{F},\frac{1+z_{a}}{D_{a}}t^{\prime}_{\mathrm{delay}}\right\}.$ (3.8)
In Eq. (3.7) in the first row we, for the sake of completeness, set $G_{\nu a}$ as zero for $t^{\gamma\max}_{a}>t_{\nu}+\Delta_{F}$, albeit technically $G_{\nu a}$ is not defined for these times because by definition $t^{\gamma\max}_{a}$ is the time of the maximum of the Fermi-LAT light curve of blazar $a$ only among times $\leq t_{\nu}+\Delta_{F}$. The second row in Eq. (3.7) assigns equal weights to all values of $t^{\gamma\max}_{a}$ if they preceed $t_{\nu}$ by $t_{a\mathrm{\,delay}}$ with an absolute margin of $\Delta_{F}$ to account for the finite size of the Fermi-LAT time bin. The third row handles all other cases assigning the Gaussian penalty with the width equal to the anticipated neutrino delay $t_{a\mathrm{\,delay}}$ in the Earth frame. The illustration of the graph of $G_{\nu a}$ as a function of $t^{\gamma\max}_{a}$ for a fixed $t_{\nu}$ and $t_{a\mathrm{\,delay}}$ is given in Fig. 1.
Figure 1: An example of the graph of the temporal part $G_{\nu a}$ of weights defined in Eq. (3.7) as a function of $t^{\gamma\max}_{a}$ for a fixed $t_{\nu}$ (shown as the red solid vertical line) and a set of fixed $t_{a\mathrm{\,delay}}$ (see the legend).
As was demonstrated by [46], the most sensitive global test statistic (TS) for the whole dataset is just the sum of individual weights:
$\mathrm{TS}(t^{\prime}_{\mathrm{delay}})=\sum\limits_{\nu}w_{\nu}(t^{\prime}_{\mathrm{delay}}),$ (3.9)
@@
We compare the obtained $\mathrm{TS}$ value for the real dataset $\{x_{\nu}\}$ against the distribution of values $\mathrm{TS}_{f}$ obtained for the set of mock datasets $\{m_{\nu}^{f}\}$ with values of right ascensions $\alpha_{\nu}^{f}$ and arrival times $t_{\nu}^{f}$ randomly generated as described in Sect. 2.1.1. Then, for each correlation search (test) performed, we, following [12, 55], obtain the p-value
$p(t^{\prime}_{\mathrm{delay}})=\frac{\#\{\mathrm{TS}_{f}\geq\mathrm{TS}\}_{f=1}^{N_{f}}+1}{N_{f}+1},$ (3.10)
@@ 4 Results
Figure 2: $p$-value (see Eq. 3.10) as a function of the jet-frame time delay assumed to be universal for all AGNs for the test with the sources from H21+ (dotted curve) and sources from R24+ (dashed-dotted curve).
@@
The results of the tests performed for H21+ and R24+ source lists are shown in Fig. 2. For the H21+ source list, the global minimum of p-value is $0.037$ reached at the jet-frame delay $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d. For the R24+ source list, the minimum of $p(\hat{t}^{\prime}_{\mathrm{delay}})=0.026$ is reached at $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d. Tables 1 and 2 present the top ten associations with the highest values of $\mathrm{TS}_{\nu}$ for H21+ and R24+ accordingly [^6] For the H21+ source list, we ran an additional test with an extra factor $F^{X}_{a}$ in Eq. (3.5), where $F^{X}_{a}$ is the relative radio flux of sources in the X radio band (between 8 GHz and 12 GHz). This test was motivated by Refs. [12, 14], that reported evidence of the correlation between the IceCube neutrino alerts and radio fluxes in the cores of radio sources at the frequency of $8$ GHz. From the RFC, we used the column with the “median flux density at X-band at baseline projection lengths shorter than 1000 km” [38] and normalised it by the maximum value among the sources in the H21+ source list. This test did not result in a significant correlation with a p-value curve similar to the one in Fig. 2..
In Fig. 2, one can see that the p-value curve of R24+ almost reaches the $2\sigma$ level (one-tailed). The trial correction (for multiple tested values of $t^{\prime}_{\mathrm{delay}}$) performed according to the prescriptions of Ref. [12] shows, however, that the minima of the pre-trial p-values as deep as in Fig. 2 occur by chance in $p_{\mathrm{post-trial\,H21+}}=14$% ($1.06\sigma$) and $p_{\mathrm{post-trial\,R24+}}=11$% ($1.21\sigma$) among mock datasets for the tests with H21+ and R24+ source lists respectively. Thus, none of the obtained p-values show evidence of a significant association between IceCat-1 neutrinos and AGNs from the considered source lists.
We note that the optimal delay $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d for H21+ is by $4$ times greater than $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d for R24+. This is driven by the substantial difference in the Doppler factors (compare Tables 1 and 2) of 3C 454.3 and 3C 279, that contribute 5353% of the total TS for the former and 8080% of the total TS for the latter: Ref. [37] estimates the Doppler factor of 3C 279 by a factor of five larger than [56], and the Doppler factor of 3C 454.3 by a factor of three larger than [56].
@@
Figure 3: Light curves of 3C 454.3 (upper panel), and 3C 279 (lower panel), which are among the most significant associations between IceCat-1 neutrinos and AGNs in both the tests (see Tables 1 and 2). The connected markers with error bars show monthly-binned Fermi-LAT photon fluxes in the energy range 100 MeV–100 GeV from the public repository [32] (ordinate axis). Dotted vertical lines indicate the time $t_{\nu}$ of the neutrino arrival. Time intervals between the alert and the global maximum of the light curve prior to $t_{\nu}$ are shown as double-pointed arrows. Vertical upper dashed risks indicate the anticipated position of the maximum with the Doppler factor from R24+ and $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d, and vertical lower dashed-dotted risks — with the Doppler factor from H21+ and $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d. The width of the Gaussian temporal factor $G_{\nu a}$ in weights is equal to the distance between the vertical risks and the line of $t_{\nu}$ (see Eq. 3.7).
The most significant association is driven by the alert IC120523B [^7] In the original publication of [25], this alert had a name IC120523A, which was also given to another alert in the catalogue; in the updated version of IceCat-1 [33], the alert has the name IC120523B. associated with the brightest Fermi-LAT blazar flare of the FSRQ 3C 454.3, which happened in November 2010 [53, 54, 27]. Despite large angular uncertainty of $47.2$ sq. deg., the exceptional brightness of 3C 454.3 during that flare and that the neutrino arrival delay in the Earth frame is not far from the anticipated $(t^{\prime}_{\mathrm{delay}}(1+z)/D)$ for the both tested source lists, provides the highest value of $\mathrm{TS}_{\nu}$ in both Tables 1 and 2. 3C 454.3 is well within the original IceCube angular uncertainty contour lying in $0.73^{\circ}$ from the best-fit arrival direction of IC120523B, and this association does not require its extension by $\Delta$.
3C 279 is the second brightest FSRQ in the Fermi-LAT 4FGL catalogue [40]. The neutrino IC150926A arrives around $100$ d (mind the finite size of one month of the Fermi-LAT light curves) after the bright but short multiwavelength flare of 3C 279 happened around MJD 57 200 detected as well in X-rays and optical photons [57, 58, 59]. The peak of the flare is in perfect agreement with the anticipated neutrino lagging by $(t^{\prime}_{\mathrm{delay}}(1+z)/D)$ for both H21+ and R24+ source lists. As we mentioned above, the four-times difference in estimates of Doppler factors by [37] and [52] seems to drive the difference in the optimal delays providing the minima of p-values in Fig. 2. The combination of the source brightness and the prior $\gamma$-flux global maximum happening at the anticipated time, together with a relatively small angular uncertainty of $8.8$ sq. deg., result in the second highest value of $\mathrm{TS}_{\nu}$ in both Tables 1 and 2. This association requires the extension of the original IceCube angular uncertainty contour by $\Delta=0.78^{\circ}$.

### v2

@@ 3 Methods
$\left(\frac{d^{\alpha}_{\nu b}}{\tilde{\alpha}_{\nu}^{\pm}}\right)^{2}+\left(\frac{d_{\nu b}^{\delta}}{\tilde{\delta}_{\nu}^{\pm}}\right)^{2}\leq 1,$ (3)
@@
$\tilde{\alpha}_{\nu}^{\pm}=\alpha_{\nu}^{\pm}+\Delta,$ (4)
$\tilde{\delta}_{\nu}^{\pm}=\delta_{\nu}^{\pm}+\Delta,$ (5)
where $d^{\alpha}_{\nu b}$ is the difference between blazar’s $b$ and neutrino’s $\nu$ RAs, $d^{\delta}_{\nu b}$ is the difference between blazar’s and neutrino’s declinations, the sign ++ or - in $\alpha_{\nu}^{\pm}$ and $\delta_{\nu}^{\pm}$ is chosen according to the sign of $d^{\alpha}_{\nu b}$ and $d^{\delta}_{\nu b}$; $\Delta=0.78^{\circ}$ is the assumed systematic angular uncertainty. This value was found (Plavin et al., 2023) to give the optimal correlation strength between the radio-bright blazars and neutrino alerts, and we keep it fixed to avoid statistical penalising for multiple trials. A similar value of $1^{\circ}$ was found (Kouch and others, 2024) to be optimal in the search for spatio-temporal correlations between the multiwavelength light curves of the CGRaBS blazars (Healey et al., 2008) and the IceCube high-energy neutrinos. In a dedicated study with synthesised catalogues and a controlled number of synthetic blazar-neutrino correlations, Kouch et al. (2025b) showed that enlarging the value of $\Delta$ does not lead to an increase in chances of finding a false-positive statistically significant correlation.
Following Eq. (1) in the work by Kouch et al. (2025b), we define for each of the alerts its enlarged angular uncertainty area as
$U_{\nu}=\frac{\pi}{4}\left(\tilde{\alpha}^{+}_{\nu}\tilde{\delta}^{+}_{\nu}+\tilde{\alpha}^{-}_{\nu}\tilde{\delta}^{+}_{\nu}+\tilde{\alpha}^{-}_{\nu}\tilde{\delta}^{-}_{\nu}+\tilde{\alpha}^{+}_{\nu}\tilde{\delta}^{-}_{\nu}\right),$ (6)
@@
If neutrinos are produced in the external photon fields, approximate linearity between the $\gamma$ and $\nu$ fluxes may be expected. According to Dermer et al. (2012), the Doppler amplification $\propto D^{p}$ makes the observed neutrino flux boosted with $p=4$ if neutrinos are produced in $p\gamma$ interactions on internal photon fields co-moving with the jet frame. However, $p\geq 5$ is expected if $p\gamma$ interactions are dominated by the photon fields external to the relativistic jet, with especially strong amplifications $p>5$ if $p\gamma$ photopion production occurs near the threshold. The $\gamma$-ray external Compton emission is, in turn, produced with amplification $\propto p^{6}$, making $F_{\nu}\propto F_{\gamma}^{\geq 5/6}\sim F_{\gamma}$, which justifies our assumption of the linear $\gamma-\nu$ flux scaling. Franckowiak and others (2020) found that neutrino-emitting blazar candidates are statistically compatible with hypotheses of both a linear correlation and no correlation between neutrino and gamma-ray energy fluxes.
We expect the jet-frame time delay $t^{\prime}_{\mathrm{delay}}$ to relate to the photopion proton energy-loss timescale $t^{\prime}_{p\gamma}$ and/or proton acceleration timescale $t^{\prime}_{\mathrm{acc}}$. They can vary significantly depending on the proton energy and parameters of photon fields serving as the target for the photopion production process, on the magnetic field, and proton acceleration mechanism. A priori, we do not know the exact value of $t^{\prime}_{\mathrm{delay}}$. Thus, we perform a scan over parameter $t^{\prime}_{\mathrm{delay}}\sim U[t^{\prime}_{\min};t^{\prime}_{\max}]$, $t^{\prime}_{\min}=\Delta_{F}=1$ month corresponding to the bin size $\Delta_{F}$ of the Fermi-LAT light curves. Podlesnyi and Oikonomou (2025), in modelling the brightest Fermi-LAT blazar flare of 3C 454.3, obtained that close to the threshold of the photopion production around $E_{\nu}\sim 100$ TeV, $t^{\prime}_{\mathrm{delay}}\approx t^{\prime}_{p\gamma}\approx t^{\prime}_{\mathrm{acc}}\sim 10^{9}$ s (see Fig. 12 by Podlesnyi and Oikonomou, 2025). Thus, we adopt $t^{\prime}_{\max}=3\times 10^{9}$ s $3.47\times 10^{4}$ d being agnostic of the exact mechanism of the possible $\nu$ delay and no particular preference to any delay between $t^{\prime}_{\min}$ and $t^{\prime}_{\max}$.
@@
$w_{\nu}(t^{\prime}_{\mathrm{delay}})=\sum_{a}\left[s_{\nu}\rho_{\nu}G_{\nu a}(t^{\gamma\max}_{a},t_{\nu};t^{\prime}_{\mathrm{\,delay}},D_{a},z_{a})F^{\gamma}_{a}(t^{\gamma\max}_{a})\right],$ (7)
where the sum $\sum_{a}$ runs over all sources within the uncertainty region of neutrino $\nu$ (as was pointed out by Kouch et al. (2025a), a single neutrino cannot come from more than one source, but having several blazars in its angular uncertainty region increases the probability of this neutrino coming from a blazar), $s_{\nu}$ is the alert signalness, $\rho_{\nu}$ is the factor penalising alerts with too large uncertainties. Following recommendations of Kouch et al. (2025b), we use a top-hat spatial factor penalising alerts with too large angular uncertainties:
$\rho_{\nu}=\begin{cases}1,\mathrm{\,if\,}U_{\nu}\leq\tilde{U}_{\nu}\\ \displaystyle{\left(\frac{\tilde{U}_{\nu}}{U_{\nu}}\right)},\mathrm{\,if\,}U_{\nu}>\tilde{U}_{\nu},\end{cases}$ (8)
where $\tilde{U}_{\nu}=15.6$ sq. deg. is the median angular uncertainty (see Eq. 6) of IceCat-1 alerts (after the extension by $\Delta=0.78^{\circ}$).
The key point of the current paper is the search for neutrino-blazar correlations taking into account the possibility of neutrino emission lagging behind the $\gamma$-ray flare due to slowness of the $p\gamma$ photopion production and/or proton acceleration as was suggested by Podlesnyi and Oikonomou (2025). Given that even during the brightest blazar flares, the IceCube neutrino yield is of the order of $10^{-3}-10^{-2}$ muon neutrinos at energies $\gtrsim 100$ TeV (Oikonomou et al., 2019; Kreter et al., 2020; Robinson and Boettcher, 2024) and during quieter activity periods even lower (assuming proportionality between $\gamma$-ray and $\nu$ fluxes), we assume that if the neutrino is indeed associated with the blazar, it must have come in a causal connection with the major flare. Thus, the last term in Eq. (7) denotes the global maximum of the Fermi-LAT light curve of blazar $a$ detected at $t^{\gamma\max}_{a}$ among the times prior to the associated neutrino arrival ($\Delta_{F}=1$ month to account for the size of the time bin and still allow for instantaneous associations) normalized by $1.66\times 10^{-5}$ cm-2 s-1 (the maximum value of monthly-averaged fluxes among the studied blazars within the whole Fermi-LAT observation history observed from FSRQ 3C 454.3 during the brightest Fermi-LAT blazar flare in November 2010 (Fermi-LAT Collaboration, 2011; Nalewajko, 2013; Podlesnyi and Oikonomou, 2025)). As for the temporal term $G_{\nu a}$ accounting for the neutrino delay, we define it linking to the time $t^{\gamma\max}_{a}$ of the global maximum of the associated blazar light curve prior to the neutrino arrival as follows
@@
$t_{a\mathrm{\,delay}}(t^{\prime}_{\mathrm{delay}},D_{a},z_{a})=\max\left\{\Delta_{F},\frac{1+z_{a}}{D_{a}}t^{\prime}_{\mathrm{delay}}\right\}.$ (10)
Figure 1: An example of the graph of the temporal part $G_{\nu a}$ of weights defined in Eq. (9) as a function of $t^{\gamma\max}_{a}$ for a fixed $t_{\nu}$ (shown as the red solid vertical line) and a set of fixed $t_{a\mathrm{\,delay}}$ (see the legend).
In Eq. (9) in the first row we, for the sake of completeness, set $G_{\nu a}$ as zero for $t^{\gamma\max}_{a}>t_{\nu}+\Delta_{F}$, albeit technically $G_{\nu a}$ is not defined for these times because by definition $t^{\gamma\max}_{a}$ is the time of the maximum of the Fermi-LAT light curve of blazar $a$ only among times $\leq t_{\nu}+\Delta_{F}$. The second row in Eq. (9) assigns equal weights to all values of $t^{\gamma\max}_{a}$ if they preceed $t_{\nu}$ by $t_{a\mathrm{\,delay}}$ with an absolute margin of $\Delta_{F}$ to account for the finite size of the Fermi-LAT time bin. The third row handles all other cases assigning the Gaussian penalty with the width equal to the anticipated neutrino delay $t_{a\mathrm{\,delay}}$ in the Earth frame. The illustration of the graph of $G_{\nu a}$ as a function of $t^{\gamma\max}_{a}$ for a fixed $t_{\nu}$ and $t_{a\mathrm{\,delay}}$ is given in Fig. 1.
As was demonstrated by Kouch et al. (2025b), the most sensitive global test statistic (TS) for the whole dataset is just the sum of individual weights:
$\mathrm{TS}(t^{\prime}_{\mathrm{delay}})=\sum\limits_{\nu}w_{\nu}(t^{\prime}_{\mathrm{delay}}),$ (11)
@@
We compare the obtained $\mathrm{TS}$ value for the real dataset $\{x_{\nu}\}$ against the distribution of values $\mathrm{TS}_{f}$ obtained for the set of mock datasets $\{m_{\nu}^{f}\}$ with values of right ascensions $\alpha_{\nu}^{f}$ and arrival times $t_{\nu}^{f}$ randomly generated as described in Sect. 2.1.1. Then, for each correlation search (test) performed, we, following Plavin et al. (2020); Davison and Hinkley (1997), obtain the p-value
$p(t^{\prime}_{\mathrm{delay}})=\frac{\#\{\mathrm{TS}_{f}\geq\mathrm{TS}\}_{f=1}^{N_{f}}+1}{N_{f}+1},$ (12)
@@ 4 Results
The results of the tests performed for H21+ and R24+ source lists are shown in Fig. 2. For the H21+ source list, the global minimum of p-value is $0.037$ reached at the jet-frame delay $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d. For the R24+ source list, the minimum of $p(\hat{t}^{\prime}_{\mathrm{delay}})=0.026$ is reached at $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d. Tables 1 and 2 present the top ten associations with the highest values of $\mathrm{TS}_{\nu}$ for H21+ and R24+ accordingly [^5] For the H21+ source list, we ran an additional test with an extra factor $F^{X}_{a}$ in Eq. (7), where $F^{X}_{a}$ is the relative radio flux of sources in the X radio band (between 8 GHz and 12 GHz). This test was motivated by Plavin et al. (2020, 2023), that reported evidence of the correlation between the IceCube neutrino alerts and radio fluxes in the cores of radio sources at the frequency of $8$ GHz. From the RFC, we used the column with the “median flux density at X-band at baseline projection lengths shorter than 1000 km” (Petrov and Kovalev, 2025) and normalised it by the maximum value among the sources in the H21+ source list. This test did not result in a significant correlation with a p-value curve similar to the one in Fig. 2..
Figure 2: $p$-value (see Eq. 12) as a function of the jet-frame time delay assumed to be universal for all AGNs for the test with the sources from H21+ (dotted curve) and sources from R24+ (dashed-dotted curve).
In Fig. 2, one can see that the p-value curve of R24+ almost reaches the $2\sigma$ level (one-tailed). The trial correction (for multiple tested values of $t^{\prime}_{\mathrm{delay}}$) performed according to the prescriptions of Plavin et al. (2020) shows, however, that the minima of the pre-trial p-values as deep as in Fig. 2 occur by chance in $p_{\mathrm{post-trial\,H21+}}=14$% ($1.06\sigma$) and $p_{\mathrm{post-trial\,R24+}}=11$% ($1.21\sigma$) among mock datasets for the tests with H21+ and R24+ source lists respectively. Thus, none of the obtained p-values show evidence of a significant association between IceCat-1 neutrinos and AGNs from the considered source lists.
We note that the optimal delay $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d for H21+ is by $4$ times greater than $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d for R24+. This is driven by the substantial difference in the Doppler factors (compare Tables 1 and 2) of 3C 454.3 and 3C 279, that contribute 5353% of the total TS for the former and 8080% of the total TS for the latter: Homan et al. (2021) estimates the Doppler factor of 3C 279 by a factor of five larger than the estimate of Rodrigues et al. (2024b), and the Doppler factor of 3C 454.3 by a factor of three larger than that found by Rodrigues et al. (2024b).
The most significant association is driven by the alert IC120523B [^6] In the original publication of Abbasi and others (2023c), this alert had a name IC120523A, which was also given to another alert in the catalogue; in the updated version of IceCat-1 (IceCube Collaboration, 2023), the alert has the name IC120523B. associated with the brightest Fermi-LAT blazar flare of the FSRQ 3C 454.3, which happened in November 2010 (Fermi-LAT Collaboration, 2011; Nalewajko, 2013; Podlesnyi and Oikonomou, 2025). Despite large angular uncertainty of $47.2$ sq. deg., the exceptional brightness of 3C 454.3 during that flare and that the neutrino arrival delay in the Earth frame is not far from the anticipated $(t^{\prime}_{\mathrm{delay}}(1+z)/D)$ for the both tested source lists, provides the highest value of $\mathrm{TS}_{\nu}$ in both Tables 1 and 2. 3C 454.3 is well within the original IceCube angular uncertainty contour lying in $0.73^{\circ}$ from the best-fit arrival direction of IC120523B, and this association does not require its extension by $\Delta$.
3C 279 is the second brightest FSRQ in the Fermi-LAT 4FGL catalogue (Abdollahi and others, 2020). The neutrino IC150926A arrives around $100$ d (mind the finite size of one month of the Fermi-LAT light curves) after the bright but short multiwavelength flare of 3C 279 happened around MJD 57 200 detected as well in X-rays and optical photons (Paliya, 2015; Pittori and others, 2018; Singh et al., 2020). The peak of the flare is in perfect agreement with the anticipated neutrino lagging by $(t^{\prime}_{\mathrm{delay}}(1+z)/D)$ for both H21+ and R24+ source lists. As we mentioned above, the four-times difference in estimates of Doppler factors by Homan et al. (2021) and Robinson and Boettcher (2024) seems to drive the difference in the optimal delays providing the minima of p-values in Fig. 2. The combination of the source brightness and the prior $\gamma$-flux global maximum happening at the anticipated time, together with a relatively small angular uncertainty of $8.8$ sq. deg., result in the second highest value of $\mathrm{TS}_{\nu}$ in both Tables 1 and 2. This association requires the extension of the original IceCube angular uncertainty contour by $\Delta=0.78^{\circ}$.
@@
Figure 3: Light curves of 3C 454.3 (upper panel), and 3C 279 (lower panel), which are among the most significant associations between IceCat-1 neutrinos and AGNs in both the tests (see Tables 1 and 2). The connected markers with error bars show monthly-binned Fermi-LAT photon fluxes in the energy range 100 MeV–100 GeV from the public repository (Abdollahi and others, 2023) (ordinate axis). Dotted vertical lines indicate the time $t_{\nu}$ of the neutrino arrival. Time intervals between the alert and the global maximum of the light curve prior to $t_{\nu}$ are shown as double-pointed arrows. Vertical upper dashed risks indicate the anticipated position of the maximum with the Doppler factor from R24+ and $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d, and vertical lower dashed-dotted risks — with the Doppler factor from H21+ and $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d. The width of the Gaussian temporal factor $G_{\nu a}$ in weights is equal to the distance between the vertical risks and the line of $t_{\nu}$ (see Eq. 9).